OpenStudy (anonymous):

Please! check my answer. A gumball machine contains 7 red, 8 orange, 9 purple, 7 white, and 5 yellow gumballs. Tyson buys 3 gumballs. Find each probability,assuming that the machine dispenses the gumballs at random. P(2 white, 1 purple) (7/36)*(6/35)*(9/34) = ?

OpenStudy (anonymous):

Looks good

OpenStudy (anonymous):

Good work man

OpenStudy (dumbcow):

since order doesn't matter, use combinations total combinations = 36 Choose 3 \[P = \frac{\left(\begin{matrix}7 \\ 2\end{matrix}\right) *\left(\begin{matrix}9 \\ 1\end{matrix}\right) }{\left(\begin{matrix}36 \\ 3\end{matrix}\right)}\]

OpenStudy (dumbcow):

your answer only gives the probability of taking a white gumball first and second then the purple one as the 3rd. But what if the purple came out first or second?

OpenStudy (dumbcow):

or did they give the order? then ignore me :)

OpenStudy (anonymous):

the question is white first, next purple

OpenStudy (dumbcow):

ok you are correct then

OpenStudy (anonymous):

my correct?

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

ty